i need help solving x^2-14x+49=9
i have gotten to x^2-14x=-40
but i dont know where to go from there

Answer by nerdybill(5411)

x^2-14x+49=9
.
The idea to move everything one side and leave zero on the other.
So, begin by subtracting 9 from both sides:
.
x^2-14x+40=0
.
Now, factor the left side:
(x-4)(x-10) = 0
.
x = {4, 10}